How do I calculate the diameter of a rubber band (circular in shape) if this rubber is to fit onto the perimeter of a rectangle 16cmx9cm. The rubber band is stretchable about 15%
Answer by ilana(305)

The perimeter of this rectangle=16+16+9+9=50cm.
Circumference=pi*Diameter.
We want piD+15%piD=50, so 115piD/100=50
115piD=5000, so 23piD=1000, so D=1000/23pi.